OpenStudy (anonymous):

WILL MEDAL (Explain as well as you can please.) The formula for the slant height of a cone is , where S is surface area of the cone. Use the formula to find the slant height, l, of a cone with a surface area of 500π ft2 and a radius of 15 ft. l =?? ft

OpenStudy (anonymous):

IF 3+3=9 I BELIEVE THE ANSWER IS B

OpenStudy (perl):

Is the formula given for the slant height ?

OpenStudy (anonymous):

here it is! Sorry!

OpenStudy (perl):

We can plug in S and r into the formula

OpenStudy (anonymous):

Yeah because those are given in the problem but aren't I suppose to solve for the variable? Or ... if I do that am I already doing that?

OpenStudy (perl):

Here we don't have to solve. just plug in :) $$ \large l=\frac{S - \pi r^2 }{\pi} $$

OpenStudy (anonymous):

Oh okay. I'll do it now.

OpenStudy (anonymous):

I got 275 ft